# Editing Knowledge Representation of Language Lodel via Rephrased Prefix Prompts

### v1

@@ 3.2.1 Concepts
Many studies show that methods based on weight-modified only perform editing on specific knowledge and cannot utilize this updated knowledge for reasoning [30]. In this section, we utilize PSPEM to control the model for effective reasoning. Specifically, we want to manipulate a person’s name or an object. As shown in Figure 2, given the prompt statements: "Tawnie enjoys indoor cycling and yoga, a good glass of red wine with friends, and of course anything to do with being in the kitchen!" and "Banana has four legs. ", the model should infer "Tawnie has the occupation of dietitian. " and "Banana is an animal. " based on its reasoning. However, traditional prompt engineering often causes the model to neglect critical information in the prompt, leading it to respond with "Tawnie has the occupation of yoga. " and "Banana is a fruit. ".
@@ 3.2.2 Datasets
We chose BioBias [6] and the McRae [18] as our foundational datasets. Biobias contains 397,000 short professional biographies of non-celebrities gathered from the internet, each marked with an occupational theme. From each biography, we extract a sentence, substituting the individual’s full name with only their first name, and use this sentence to prompt the language model (LM) by appending "{Person} has the occupation of...", as shown in the middle example on the left in Figure 1. We subsequently assess the language model’s accuracy by examining the relative probabilities assigned to 28 potential occupations, deeming the model correct if it ranks the individual’s actual occupation as the most likely. McRae encompasses 541 concepts, 2,526 features, and details on how frequently each feature was identified as prototypical for each concept by human evaluators. [13]. Following Hernandez et al. [13], we construct a dataset comprising 10,000 entries. Each entry comprises a concept $c$, a list of original features $f^{(o)}$ for the concept, a target feature to be added $f^{*}$, and a list of features $f^{(c)}$ that correlate with the new feature. For example, we use the common noun "Banana" as the editing target and the feature description "has four legs" as the attribute. Properties such as "animal" exist in a complex network of entailment and correlation relations. We hope that based on the prompt information, LMs can respect these relations (e.g., given a prompt "Banana has four legs", LMs can increase the probability that "Banana is an animal" and decrease the probability that "Banana cannot move freely").
@@ 3.2.4 Baselines
    REMEDI works by extracting attribute information from a prompt and then injecting this information into the subject word via a linear transformation.
@@ 3.2.6 Results
Table 3 shows the evaluation results on BioBias. Without prompts, the model exhibits random guesses of the answers due to the inclusion of non-celebrity names in BioBias, as the model itself does not store information about these names. Subsequently, it generated non-repetitive but unordered text. By guiding the model with prompts, the model’s estimation of occupations improved, reaching 57.9% and 55.5% respectively, but still lagging behind trained methods in terms of consistency in text generation. While REMEDI shows improvement in accuracy, this is at the expense of the quality of the generated text, as low Fluency and Consistency indicate that REMEDI consistently produces repetitive and irrelevant text. PSPEM, on the other hand, demonstrated robust performance, not only achieving the highest prediction accuracy but also ensuring the fluency and consistency of the generated text.
@@
Table 4 shows the evaluation results on McRae. We evaluate the effects of different methods from two aspects: by providing an object with a conceptual prompt $c$, such as "Banana has four legs. ", “Original” evaluates the degree of change in the object on the original features $f^{(o)}$, such as "Banana cannot move freely". “Related”, evaluates the degree of change for features $f^{(c)}$ related to the target feature $f^{*}$, such as "Banana is an animal.". Without prompts, the model can identify the original concepts of the objects. After adding conceptual prompts, although the model can associate the object with $f^{(c)}$, it is unable to forget $f^{(o)}$. In terms of associating $f^{(c)}$, REMEDI and PSPEM perform better. Compared with REMEDI, while the performance effect of PSPEM was lower than REMEDI for associating the relevant features, it was more substantial than REMEDI for forgetting the original features, and it’s worth mentioning that rather than forgetting these features, REMEDI enhances the association with $f^{(o)}$. It implies that REMEDI has not effectively learned and incorporated the prompt information.
@@ 3.4 Ablation Study
    We explored how we could better extract information from the prompts, specifically, we tried to use the methods in REMEDI [13] to extract attribute information instead of extracting information from the entire sentence.
@@
Table 5 shows the results of ablation experiments. We observe that the information extraction method proposed by REMEDI performs poorly, showing lower performance compared to ours in almost all hyperparameter settings. This indicates that PSPEM better extracts information from prompt sentences. On the other hand, except for the hyperparameter choices of $\lambda_{1}$=1 and $\lambda_{2}$=1, more or less failures are observed in other settings. Hyper-parameter $\lambda_{1}$ controls the accuracy of editing or prediction, with higher proportions of $\lambda_{1}$ leading to higher accuracy. Hyper-parameter $\lambda_{2}$ controls the quality of the generated text, with models having lower $\lambda_{2}$ often yielding poorer results in metrics such as GS, RS, Flu, or Con. The best performance is achieved only when $\lambda_{1}$=1 and $\lambda_{2}$=1.
Table 5: The results of the ablation experiments on GPT-J-6B, “Attr” denotes the use of the information extraction method in REMEDI and “Ours” denotes the use of the method proposed in this paper. Red numbers indicate poor metrics.
@@ 3.5 Similarity To Prompt
$*\displaystyle D_{i}^{l}*$ $=\text{softmax}(h_{i}^{l}\text{(PROMPT)}\cdot W_{E}),$
@@
    RAND. We evaluate using randomly guessed answers or randomly generated vectors of the same dimension.
@@
Table 6 summarizes the similarities between each method and the PREFIX PROMPT in the COUNTERFACT and Biobias datasets using GPT2-XL. For metrics that require measuring the internal performance of the model, we calculate the final evaluation value by computing the average of the last five layers of the model. The results indicate that PSPEM performed the best across these four metrics. Moreover, both the CosSim and KL metrics suggest that the information encoded by PSPEM has a high resemblance to the prompt, revealing why PSPEM can accurately edit model outputs based on the original prompt information without deviating from it.
@@ 4 CONCLUSION
This paper introduces a prompt-based knowledge editing approach, PSPEM. PSPEM employs a two-step process, involving compression and refinement, to extract crucial information from prompts. It uses alignment techniques to guide the model in generating text that remains consistent with the given prompts. PSPEM integrates the advantages of two types of editing methods, and showcases robust performance in both knowledge editing and attribute inserting. It not only enables the editing of individual pieces of information but also allows for reasoning and the modification of concepts of existing objects based on the provided prompts. Similarity experiments highlight that one of PSPEM’s outstanding strengths lies in its ability to emulate the influence of original prompts on the model’s internal processes. We believe our findings will promote the application of prompt engineering in the field of knowledge editing and catalyze the development of more user-friendly and accurate LM editing tools.

### v2

@@ 3.2.1 Concepts
Many studies show that methods based on weight-modified only perform editing on specific knowledge and cannot utilize this updated knowledge for reasoning [30]. To address this, we apply PSPEM for more effective model reasoning, especially in manipulating complex concepts such as personal names or objects in non-traditional contexts. As shown in Figure 1, given the prompt statements: "Tawnie enjoys indoor cycling and yoga, a good glass of red wine with friends, and of course anything to do with being in the kitchen!" and "Banana has four legs. ", the model should infer "Tawnie has the occupation of dietitian. " and "Banana is an animal. " based on its reasoning. However, traditional prompt engineering often causes the model to neglect critical information in the prompt, leading it to respond with "Tawnie has the occupation of yoga. " and "Banana is a fruit. ". These scenarios demonstrate PSPEM’s ability to guide the model through complex reasoning from deliberately misleading or non-standard information.
However, it’s important to note that such examples are designed to test the limits of PSPEM’s reasoning capabilities, especially in contrast to traditional prompt engineering methods, which might lead to oversimplified or incorrect conclusions like "Tawnie has the occupation of yoga. " and "Banana is a fruit. " due to their inability to adequately prioritize or analyze the given prompt information.”
@@ 3.2.2 Datasets
We chose BioBias [6] and the McRae [18] as our foundational datasets. Biobias contains 397,000 short professional biographies of non-celebrities gathered from the internet, each marked with an occupational theme. From each biography, we extract a sentence, substituting the individual’s full name with only their first name, and use this sentence to prompt the language model (LM) by appending "{Person} has the occupation of...", as shown in the middle example on the left in Figure 1. We subsequently assess the language model’s accuracy by examining the relative probabilities assigned to 28 potential occupations, deeming the model correct if it ranks the individual’s actual occupation as the most likely. McRae encompasses 541 concepts, and 2,526 features, and details on how frequently each feature was identified as prototypical for each concept by human evaluators. [13]. Following Hernandez et al. [13], we construct a dataset comprising 10,000 entries. Each entry comprises a concept $c$, a list of original features $f^{(o)}$ for the concept, a target feature to be added $f^{*}$, and a list of features $f^{(c)}$ that related with the new feature. For example, we use the common noun "Banana" as the editing target and the feature description "has four legs" as the attribute. Properties such as "animal" exist in a complex network of entailment and correlation relations. We hope that based on the prompt information, LMs can respect these relations (e.g., given a prompt "Banana has four legs", LMs can increase the probability that "Banana is an animal" and decrease the probability that "Banana cannot move freely").
@@ 3.2.4 Baselines
    REMEDI works by extracting attribute information from a biography prompt and then injecting this information into the subject word via a linear transformation.
@@ 3.2.6 Results
Table 3 shows the evaluation results on BioBias. Without prompts, the model exhibits random guesses of the answers due to the inclusion of non-celebrity names in BioBias, as the model itself does not store information about these names. Subsequently, it generated non-repetitive but unordered text. By guiding the model with biography prompts, the model’s estimation of occupations improved, reaching 57.9% and 55.5% respectively, but still lagging behind trained methods in terms of consistency in text generation. While REMEDI shows improvement in accuracy, this is at the expense of the quality of the generated text, as low Fluency and Consistency indicate that REMEDI consistently produces repetitive and irrelevant text. PSPEM, on the other hand, demonstrated robust performance, not only achieving the highest prediction accuracy but also ensuring the fluency and consistency of the generated text.
@@
Table 4 shows the evaluation results on McRae. We evaluate the effects of different methods from two aspects: by providing an object with a conceptual prompt $c$, such as "Banana has four legs. ", “Original” evaluates the degree of change in the object on the original features $f^{(o)}$, such as "Banana cannot move freely". “Related”, evaluates the degree of change for features $f^{(c)}$ related to the target feature $f^{*}$, such as "Banana is an animal.". Without prompts, the model can identify the original concepts of the objects. After adding conceptual prompts, although the model can associate the object with $f^{(c)}$, it is unable to forget $f^{(o)}$. In terms of associating $f^{(c)}$, REMEDI and PSPEM perform better. Compared with REMEDI, while the performance effect of PSPEM was lower than REMEDI for associating the related features, it was more effective than REMEDI for forgetting the original features, and it’s worth mentioning that rather than forgetting these features, REMEDI enhances the association with $f^{(o)}$. It implies that REMEDI has not effectively learned and incorporated the prompt information.
@@ 3.4 Ablation Study
    We explored how we could better extract information from the prompts, specifically, we tried to use the methods in REMEDI [13] to extract attribute information instead of extracting information from the entire sentence. As shown in Figure 1, REMEDI attempts to extract information from "born in America" rather than "Danielle Darrieux was born in America". Please note that REMEDI differs from PSPEM not only in this aspect. What we are discussing here is which method is more effective for extracting information from the prompts.
@@
Table 5 shows the results of ablation experiments. We observe that the information extraction method proposed by REMEDI performs poorly, showing lower performance compared to ours in almost all hyper-parameter settings. This indicates that PSPEM better extracts information from prompt sentences. On the other hand, except for the hyperparameter choices of $\lambda_{1}$=1 and $\lambda_{2}$=1, more or less failures are observed in other settings. Hyper-parameter $\lambda_{1}$ controls the accuracy of editing or prediction, with higher $\lambda_{1}$ leading to higher accuracy. Hyper-parameter $\lambda_{2}$ controls the quality of the generated text, with models having lower $\lambda_{2}$ often yielding poorer results in metrics such as GS, RS, Flu, and Con. The best performance is achieved only when $\lambda_{1}$=1 and $\lambda_{2}$=1.
Table 5: The results of the ablation experiments on GPT-J-6B, “Attr” denotes the information extraction method in REMEDI, and “Ours” denotes the method proposed in this paper. Red numbers indicate poor results.
@@ 3.5 Similarity To Prompt
$D_{i}^{l}*$ $=\text{softmax}(h_{i}^{l}\text{(PROMPT)}\cdot W_{E}),$
@@
    RAND, we evaluate using randomly guessed answers or randomly generated vectors of the same dimension.
@@
Table 6 summarizes the similarities between each method and the PREFIX PROMPT in the COUNTERFACT and Biobias datasets using GPT2-XL. For metrics that require measuring the internal performance of the model, we calculate the final evaluation value by computing the average of the last five layers of the model. The results indicate that PSPEM performed the best across these four metrics. $CosSim$ results are close to 1, indicating a high degree of attention similarity between PSPEM and the original prompt in the last five layers. Furthermore, the Kullback-Leibler (KL) divergence metric being close to 0 further underscores the minimal discrepancy in the model output distribution between PSPEM and the original prompt. These findings robustly validate PSPEM’s efficacy in accurately editing model outputs and aligning with original prompt information without deviation.
These results not only showcase PSPEM’s advantages in maintaining similarity with the original prompts but also underscore its potential application in tasks involving knowledge editing and attribute inserting. By fine tuning model outputs to match specific prompt information, PSPEM offers a reliable methodology for efficiently and accurately editing language models.
@@ 4 CONCLUSION
This paper presents PSPEM, an innovative prompt-based knowledge editing method, which seamlessly integrates a two-step process of compression and refinement to accurately extract and utilize crucial information from prompts. By employing alignment techniques, PSPEM ensures the generated text remains in harmony with the intended prompts, combining the high editing success of the weight-modified method with the ability to reason from given knowledge. Our experiments demonstrate PSPEM’s robust performance in knowledge editing and attribute inserting tasks, notably highlighting its extraordinary advantage in imitating the influence of original prompts on the internal of the model.
The findings underscore PSPEM’s potential to significantly advance the application of prompt engineering in knowledge editing, setting the stage for the evolution of more intuitive and precise language model (LM) editing tools. By facilitating a deeper alignment between model outputs and human-intended meanings, PSPEM not only enhances the accuracy of knowledge representation within LMs but also broadens the scope for their application across diverse domains. We anticipate that our contributions will act as a catalyst for further research in this area, ultimately leading to the development of more user-friendly and accurate LM editing tools.